# Conductivity of concentrated salt solutions

### v4

@@ III Results and discussion
The calculation from Eq.(14) is also included in Fig. 3. It can be seen that deviations from Eq. (13) are extremely small. The difference between (13) and (14) is always below a few %. We thus conclude that compact approximate Eq. (14), which is easy to handle, can safely be employed to interpret data for inorganic salts or for a predictive purpose.
It follows from (5) that $\varrho_{h}\simeq 3.379R_{h}\sqrt{c_{\infty}}$ for a fixed $T=298$ K, where $R_{h}$ should be taken in nm. This implies that the values of $K/K_{0}$ for any univalent salt solution plotted against (dimensionless) $R_{h}\sqrt{c_{\infty}}$ should collapse into a single curve. That this is indeed so is demonstrated in Fig. 4, where we plot in a semi-log scale a relative conductivity calculated from (14) vs. $R_{h}\sqrt{c_{\infty}}$ along with the experimental data for a variety of inorganic salts at $T=298$[27, 29, 30, 31, 28]. Besides salts presented in Table 1 we include KCl, LiCl, and LiClO4 using their $R_{h}$ = 0.127, 0.163, and 0.174 nm, correspondingly, to re-scale the concentration [22]. Note that since the theory is justified provided $R_{h}\sqrt{c_{\infty}}$ is below ca. 0.305, we employ for this plot only data for $c_{\infty}\leq 3$ mol/l. One can see that the approximate theory is in a good agreement with the data for all salts, perhaps with only insignificant discrepancy at low dilution that is probably comparable with the experimental error. The discrepancy is always in the direction of smaller $K/K_{0}$ than predicted by Eq. (14), but the data for KBr and KCl show slightly larger $K/K_{0}$. The reason for qualitative differences only for these two salts is unclear and is open for discussion, but it must be remembered that this was a first-order calculation only, and we did not expect it to be very accurate. The curve calculated from Eq. (15) is also included in Fig. 4. Clearly, the data obtained at high salt are irreconcilable with this linear equation.
@@
To examine the possible discrepancy from data more closely, the results for salts from Fig. 4 are reproduced in Fig. 5, where a decrement to a relative conductivity, $1-K/K_{0}$, is plotted against $R_{h}\sqrt{c_{\infty}}$ in a log-log scale. It can be seen that in dilute solutions, where the decrease in relative conductivity is so small that was even not discernible on the scale of Fig. 4, the data are slightly above the theory indicating some additional retardation in the ion migration. An explanation can be obtained if we invoke the relaxation effect, which is due to an emergence of an additional (retarding) field caused by a distortion of the EDL symmetry during the migration. Onsager showed that when $\varrho_{h}\ll 1$ the electrophoretic and relaxation terms are summed up [9], which in our notations yields
@@
Indeed, Eq. (16) provides an excellent fit to the data up to $R_{h}\sqrt{c_{\infty}}\simeq 6\times 10^{-3}$ (millimolar concentrations). Upon increasing concentration further the experimental data begin to approach to calculations from Eq. (14), and in concentrated solutions, $R_{h}\sqrt{c_{\infty}}\geq 8\times 10^{-2}$ (or $c_{\infty}\geq 0.5$ mol/l), our theory that includes an electrophoretic retardation solely is in excellent agreement with experiment. It is natural to conclude that when dilution is low the relaxation effect reduces with $c_{\infty}$ and disappears at high concentrations. Experimental observations [32, 33], simulations [6], and (numerical) calculations based on a nonlocal electrostatics approach [18] produce an endorsement of our conclusion. We also recall, that Overbeek [34] argued that, although in highly dilute solutions the relaxation effect grows with $c_{\infty}$ (see Eq. (16)), after reaching a maximum this correction to electrophoresis of colloid particles should reduce and disappears at high salt. Exactly the same was found here for inorganic salt solutions, which supports the concept of point-particle duality of ions.
@@ IV Concluding remarks
  • The electrophoresis of inorganic ions can be accommodated within a theoretical framework that employs only a harmonic mean of hydrodynamic radii of ions and relies on a modified mean-field description of electrostatics (a concept of point-particle duality) and on a classical approach to colloid electrophoresis. The need to invoke specific constants as (arbitrary) parameters, or to correct an ionic concentration is thereby removed;
  • The conductivity of concentrated salt solutions is dictated solely by an ionic electrophoresis. Although the electrophoretic retardation dominates in highly dilute solutions too, to arrive at a more accurate description of their (in fact, extremely small) conductivity reduction, it seems necessary to include the relaxation correction.
This bears on the whole question of what we mean by a degree of dissociation that is often inferred from the conductivity measurements [36]. Our results show that the conductivity is interpreted without invoking the formation of ionic pairs, thus supporting the notion of complete dissociation of strong electrolytes. However remote from mainstream thinking this conclusion may seem, it would be worthwhile to recall that there are still lingering doubts about the reality of ion pairing, at least for univalent electrolytes in high permittivity solvents [5, 37].
@@ Acknowledgements.
This research was supported by the Ministry of Science and Higher Education of the Russian Federation. We are indebted to E .S. Asmolov, R. Buchner, Y. A. Budkov, G. T. Hefter, B. Rotenberg, and G. A. Tsirlina for feedback and advises.

### v5

@@ III Results and discussion
The calculation from Eq. (18) is also included in Fig. 5. It can be seen that deviations from Eq. (17) are extremely small. The difference between (17) and (18) is always below a few percent. We thus conclude that compact approximate Eq. (18), which is easy to handle, can safely be employed to interpret data for inorganic salts or for a predictive purpose.
It follows from (5) that $\varrho_{h}\simeq 3.379R_{h}\sqrt{c_{\infty}}$ for a fixed $T=298$ K, where $R_{h}$ should be taken in nm. This implies that the values of $K/K_{0}$ for any univalent salt solution plotted against (dimensionless) $R_{h}\sqrt{c_{\infty}}$ should collapse into a single curve. That this is indeed so is demonstrated in Fig. 6, where we plot in a semi-log scale a relative conductivity calculated from (18) vs. $R_{h}\sqrt{c_{\infty}}$ along with the experimental data for a variety of inorganic salts at $T=298\leavevmode\nobreak$[33, 35, 36, 37, 34]. Besides salts presented in Table 1 we include KCl, LiCl, and LiClO4 using their $R_{h}$ = 0.127, 0.163, and 0.174 nm, correspondingly, to re-scale the concentration [26]. Note that we employ for this plot data for $c_{\infty}$ up to 4 mol/l. One can see that the approximate theory is in a good agreement with the data for all salts, perhaps with only insignificant discrepancy at intermediate dilution. The discrepancy is always in the direction of slightly smaller $K/K_{0}$ than predicted by Eq. (18). However, at low dilution our theoretical calculations practically coincide with the experimental results. Only the data for KBr and KCl show slightly larger $K/K_{0}$. Note that there exist some data [38] suggesting that conductivity of potassium salts is even higher than in specimen examples used here, but not supported by other experiments. The reason for qualitative differences only for these two salts is unclear and is open for discussion, but it must be remembered that this was a first-order calculation only, and we did not expect it to be very accurate. The curve calculated from Eq. (19) is also included in Fig. 6. Clearly, the data obtained at high salt are irreconcilable with this linear equation.
@@
To examine the possible discrepancy from data more closely, the results for salts from Fig. 6 are reproduced in Fig. 7, where a decrement to a relative conductivity, $1-K/K_{0}$, is plotted against $R_{h}\sqrt{c_{\infty}}$ in a log-log scale. It can be seen that in dilute solutions, where the decrease in relative conductivity is so small that was even not discernible on the scale of Fig. 6, the data are slightly above the theory indicating some additional retardation in the ion migration. The discrepancy looks constant at smaller values $R_{h}\sqrt{c_{\infty}}$, which indicates an increase in the extra retardation effect with salt, but then the deviations from our (electrophoretic) theory reduce and completely disappear in concentrated solutions.
An explanation can be obtained if we invoke the relaxation effect. Onsager showed that when $\varrho_{h}\ll 1$ the electrophoretic and relaxation terms can be considered separately and are summed up [9], which in our notations yields
@@
Indeed, Eq. (20) provides an excellent fit to the data up to $R_{h}\sqrt{c_{\infty}}\simeq 6\times 10^{-3}$ (millimolar concentrations) as displayed in Fig. 7.
Upon increasing concentration further the experimental data begin to approach the results of calculations from Eq. (18), and in concentrated solutions, $R_{h}\sqrt{c_{\infty}}\geq 8\times 10^{-2}$ (or $c_{\infty}\geq 0.5$ mol/l), the theory that includes an electrophoresis solely is in excellent agreement with experiment. It is natural to conclude that, although in highly dilute solutions the relaxation effect grows with $c_{\infty}$ [see Eq. (20)], after reaching a maximum this correction to electrophoresis of inorganic salt ions should reduce and disappears at high salt. As the side note, exactly the same trends have been predicted already by Overbeek [28] for standard colloid particles.
Experimental observations [39, 40], integral equations [6], and (numerical) calculations based on a nonlocal electrostatics approach [18] support our conclusion about disappearance of relaxation at high salt. In any event, it seems even intuitively clear that the relaxation, which is caused by an emergence of an additional retarding field due to a distortion of the cloud symmetry, should vanish at very high salt: no EDLs - nothing to distort, and (as a second order effect) is always smaller than electrophoretic effect. However, the magnitude of the relaxation term in the approximate conductivity equation derived by Avni et al. [14] increases monotonically with salt, as highlighted in recent publications [6, 20]. When $a=O(\lambda_{D})$ this term becomes of the same order as electrophoretic, and in more concentrated solutions even dominates. As a result, when the EDL nearly disappears its relaxation grows. Clearly, it cannot be so, and in essence, the equation by Avni et al. [14] does not apply at molar concentrations [41]. However, simple conductivity equation (18), which provides a physically correct asymptotic behavior of the electrophoretic term and neglects relaxation, is very well suited to describe concentrated salt solutions.
@@ IV Concluding remarks
  • The electrophoresis of inorganic ions can be accommodated within a theoretical framework that employs only a harmonic mean of hydrodynamic radii of ions and relies on a mean-field description of electrostatics and on a classical approach to colloid electrophoresis. The need to invoke specific constants as (arbitrary) parameters, or to correct an ionic concentration is thereby removed;
  • The conductivity of concentrated salt solutions is dictated solely by ionic electrophoresis. Although the electrophoretic retardation dominates in highly dilute solutions too, to arrive at a more accurate description of their (in fact, extremely small) conductivity reduction, it seems necessary to include the relaxation correction.
This bears on the whole question of what we mean by a degree of dissociation that is often inferred by chemists from the conductivity measurements [42]. Our results show that the conductivity is interpreted without invoking the formation of ionic pairs, thus supporting the notion of complete dissociation of strong electrolytes. However remote from mainstream thinking this conclusion may seem, it would be worthwhile to recall that there are still lingering doubts about the reality of ion pairing, at least for univalent electrolytes in high permittivity solvents [5, 24, 43].
@@ Acknowledgements.
This research was supported by the Ministry of Science and Higher Education of the Russian Federation. We are indebted to E. S. Asmolov, R. Buchner, G. T. Hefter, B. Rotenberg, and G. A. Tsirlina for feedback and advice.
DATA AVAILABILITY
The data that support the findings of this study are available within the article.
AUTHOR DECLARATIONS
The authors have no conflicts to disclose.